# Representing local protein environments with atomistic foundation models

### v2

@@ 7 Interpreting AFM representations
While atomistic features from AFMs excel at representing local protein structure and chemistry, it remains unclear how their internal representations respond to structural perturbations or whether they encode meaningful physical properties. To address this, the behavior of AFM representations under specific conformational changes is examined. Such changes induce smooth, interpretable trajectories in the embedding space, suggesting that the features capture physical motions and respect constraints such as locality, continuity, and symmetry. These properties are explored in the following case studies.
@@
Rotated phenylalanine sidechain aromatic ring. Using the synthetically modified phenylalanine described above, we perform PCA on the MACE embeddings obtained at every $\chi_{2}$ angle for the nearby CA atoms (residues $50$ and $51$ of PDB ID: 1ZV6). We observe that the first two principal components trace a smooth, one-dimensional curve, capturing the expected $180^{\circ}$ periodicity of the rotation. However, the effect diminishes with distance: residue $49$ begins to deviate from the periodic pattern, while distant residue $40$ exhibits no discernible structure, illustrating the spatial locality of the descriptor’s sensitivity.
Helix unfolding. In context of MD simulations of helix-to-strand transition described earlier (Fig. A.10), we perform PCA on the MACE embeddings of CA atoms in residues $6$ and $7$, and obseve that they follow smooth, non-intersecting trajectories. These trajectories reflect the gradual unfolding process, indicating that the descriptor space encodes continuous and interpretable structural transitions.
Further exploration of reconstructing protein environments from AFM embeddings (inverse embedding) is presented in Appendix. I.1.
8 Related Work
Recent work has highlighted the versatility of pretrained AFMs as learned representations for molecular tasks. shiota2024universal use MACE and M3GNet features to predict chemical shifts in small molecules, while elijovsius2024zero employ MACE embeddings for zero-shot molecular generation via evolutionary search. Unlike these small-molecule applications, our method is specifically designed for proteins — large, structurally diverse biomolecules requiring environment-specific representations to capture their complex local chemistry. Similar to our idea, gokcan2022prediction predict protein pKa values using ANI-2x AFM representations; as shown in Table B.3, our approach outperforms theirs in pKa prediction accuracy.
@@ 9 Discussion
In this work, we introduced a new representation for local environments in proteins, and demonstrated its effectiveness across applications. While our downstream models used frozen AFM embeddings, task-specific fine-tuning of AFMs is a promising direction that may yield further gains. Moreover, our chemical shift prediction for HA, while accurate, still slightly lags behind UCBShift, pointing to opportunities for future improvement. Notwithstanding these limitations, we believe that the proposed fully differentiable chemical shift predictor can be used to guide AlphaFold and similar generative models maddipatla2025inverseproblemsexperimentguidedalphafold for determining structure from experimentally measured chemical shifts – an important task in protein NMR that has so far been hindered by the complexity of structure-to-chemical-shift relation.
Societal impact. The work presented here attempts to advance protein representation, since proteins are fundamental ingredients of life and are implicated in health and disease, the potential impact on basic and applied research may be profound. We do not see any special ethical concerns worth highlighting.
Appendix A Additional Figures
@@
Figure A.7: Structure of the MACE embedding space for the rotating phenylalanine aromatic ring from Fig. 2. Left: a fragment of a protein with simulated rotation of the phenylalanine sidechain ring. Right: Mace embeddings of backbone CA atoms projected onto the first two principal components. The vertical axis separates the different rotation angles for visual clarity. Note that residues close to the ring ($50$ and $51$) exhibit a one-dimensional structure with nearly perfect $180$-degree periodicity. The farther residue $49$ shows some breakdown of this structure, while residue $40$, uninfluenced by the ring current, manifests a lack of relation between ring rotation and the embedding space structure.
Figure A.8: MACE embedding space reveals meaningful structural and chemical information. Depicted are two-dimensional UMAP coordinates of $165,913$ protein environments from $1327$ non-redundant chains predicted by AlphaFold2 [29], labeled left-to-right, top-to-bottom according to the DSSP secondary structure class (Table B.6), amino acid chemical identity, the pair of backbone dihedral angles $(\phi,\psi)$, and CA secondary chemical shift (relative to a random coil).
@@
Figure A.10: Structure of the MACE embedding space for the unfolding helix from Fig. A.6. Depicted are Mace embeddings of the backbone CA atoms of residues $6$ and $7$ projected onto the first two principal components. Observe an essentially one-dimensional structure of the trajectory in the PCA space.
Figure A.11: Distribution of the secondary chemical shifts of the two stable conformers of 4OLE. Left: $100$ ns MD simulations [57] of the two conformers (marked as A (containing a helix) and B (containing a linearly structured loop) according to the original PDB annotation). Lysine 63 is highlighted. Right: secondary shift distributions of lysine 63 CA (top) and CB (bottom) atoms over the MD trajectory. Note that the helical conformer exhibits an expected excess in the CA secondary shift and a defect in the CB secondary shift.
Figure A.12: Similarity of local protein environments using MACE and LoCoHD. In each plot, environments belonging to different class labels are represented as points while pairwise Euclidean distances approximate the dissimilarities as measured using the LoCoHD metric [18] and a metric between MACE likelihoods. Circles indicate each class variability as seen by the metric. Shown is conditioning by secondary structure (top row) and amino acid identity (bottom row). We conclude that MACE captures much better the similarities between related chemical and structural classes, even if LoCoHD has been explicitly designed for these tasks.
Figure A.13: Optimized structure prediction using MACE descriptors. Shown in green is the structure of the protein 4OLE:B as predicted by AlphaFold3. The structure has an alternative conformation in the region $60-68$, (shown in red is the crystallographic structure from the PDB), whose MACE environment descriptors were used to optimize the AlphaFold prediction. The solution of the optimization problem is depicted in pink.
Appendix B Quantitative Results & Tables
@@
AFM-Features + GCN [34] (Ours)
@@
Metric Class LocoHD Egret MACE OrbNet AIMNet
Precision [%] Weighted Mean $76.947$ $95.685$ $\mathbf{95.694}$ $92.500$ $94.716$
Mean $76.424$ $\mathbf{95.540}$ $95.528$ $92.233$ $94.522$
$\alpha$-helix $80.493$ $96.466$ $\mathbf{96.657}$ $93.900$ $95.762$
$\beta$-strand $75.623$ $\mathbf{96.248}$ $95.969$ $93.749$ $95.494$
Other $73.158$ $93.905$ $\mathbf{93.958}$ $89.049$ $92.309$
Recall [%] Weighted Mean $77.334$ $\mathbf{98.636}$ $95.717$ $92.574$ $94.752$
Mean $75.532$ $\mathbf{98.508}$ $95.355$ $91.977$ $94.335$
$\alpha$-helix $87.227$ $\mathbf{97.546}$ $97.522$ $95.548$ $96.805$
$\beta$-strand $89.035$ $\mathbf{97.967}$ $97.947$ $96.276$ $97.481$
Other $50.333$ $90.512$ $\mathbf{90.597}$ $84.106$ $88.720$
F1 [%] Weighted Mean $76.323$ $\mathbf{98.633}$ $95.693$ $92.509$ $94.720$
Mean $75.048$ $\mathbf{98.552}$ $95.427$ $92.073$ $94.412$
$\alpha$-helix $83.725$ $97.003$ $\mathbf{97.087}$ $94.717$ $96.281$
$\beta$-strand $81.783$ $\mathbf{97.100}$ $96.948$ $94.996$ $96.477$
Other $59.636$ $92.177$ $\mathbf{92.247}$ $86.507$ $90.479$
@@
Table B.2: Amino acid prediction accuracy. Reported are precision, recall, and F1 scores, each calculated with respect to the ground truth amino acid class. The best-performing model for each metric is indicated in bold. Unlike secondary structure prediction, per-class performance is not reported here due to the large number ($20$) of amino acid types.
AFM-Features + GCN [34] (Ours)
@@
Metric Class LocoHD Egret MACE OrbNet AIMNet
Precision [%] Weighted Mean $97.794$ $\mathbf{99.129}$ $98.448$ $95.967$ $98.665$
Mean $98.012$ $\mathbf{98.758}$ $98.031$ $95.090$ $98.407$
Recall [%] Weighted Mean $97.805$ $\mathbf{99.133}$ $98.431$ $95.955$ $98.664$
Mean $97.699$ $\mathbf{98.932}$ $98.262$ $95.072$ $98.457$
F1 [%] Weighted Mean $97.785$ $\mathbf{99.130}$ $98.435$ $95.959$ $98.663$
Mean $97.839$ $\mathbf{98.842}$ $98.137$ $95.078$ $98.431$
Table B.3: Acid Dissociation constant (pKa) prediction accuracy. Reported are mean absolute errors (MAE) relative to the pKa values estimated using the Poisson-Boltzmann solver PypKa [55] on four amino acids. The best-performing model for each metric is indicated in bold.
AFM-Features + GCN [34] (Ours)
Residues # Samples PropKa pKa-ANI MACE OrbNet AIMNet
Glutamic acid $12592$ $0.551$ $0.445$ $0.306$ $0.306$ $\mathbf{0.265}$
Aspartic acid $9841$ $0.469$ $0.473$ $0.280$ $0.284$ $\mathbf{0.267}$
Lysine $12009$ $0.393$ $0.401$ $0.320$ $0.282$ $\mathbf{0.270}$
Histidine $1182$ $0.561$ $0.426$ $0.424$ $0.441$ $\mathbf{0.380}$
@@
Table B.4: Chemical shift prediction accuracy. Reported is the mean absolute error (MAE), in ppm, relative to the experimentally measured chemical shift values from RefDB [69] for different atoms. The best-performing model for each metric is indicated in bold.
AFM-features + GCN [34] (Ours)
@@
Atom UCB Shift LocoHD ESM MACE Egret Orb AIMNet
HA $\mathbf{0.165}$ N/A $0.247$ $0.180$ $0.176$ $0.200$ $0.201$
H $0.324$ N/A $0.402$ $\mathbf{0.300}$ $0.295$ $0.333$ $0.325$
CA $0.653$ $1.333$ $0.872$ $\mathbf{0.584}$ $0.599$ $0.683$ $0.660$
N $1.891$ $2.877$ $2.455$ $\mathbf{1.642}$ $1.667$ $2.015$ $1.875$
CB $0.758$ $1.140$ $0.923$ $0.716$ $\mathbf{0.705}$ $0.792$ $0.780$
C $0.827$ $1.297$ $1.011$ $0.744$ $\mathbf{0.743}$ $0.847$ $0.824$
CG $0.595$ $0.647$ $0.690$ $0.567$ $\mathbf{0.563}$ $0.638$ $0.611$
CD $0.615$ $0.512$ $0.550$ $0.456$ $\mathbf{0.448}$ $0.521$ $0.482$
CD2 $1.116$ $1.195$ $1.335$ $1.035$ $1.085$ $1.201$ $1.749$
CG2 $0.773$ $0.833$ $0.927$ $\mathbf{0.713}$ $0.720$ $0.760$ $0.742$
CE $0.519$ $2.295$ $0.568$ $0.447$ $\mathbf{0.436}$ $0.451$ $0.446$
@@
Table B.5: Atomistic Foundation Model (AFM) metadata. Listed are the model type, dimensionality, geometric properties of the representations, and training dataset for AFMs in Section 2.
@@
MACE large $448$ $S_{n}$ $O(3)$ invariant QM9 [54], MD22 [11], ANI-1x [64]
OrbNet-v2 $256$ $S_{n}$ $E(3)$ invariant QM9, GDB13 [10], DrugBank [39]
AIMNet2 $256$ $S_{n}$ $E(3)$ invariant ANI-1x
Egret1 $384$ $S_{n}$ $O(3)$ invariant QM9, MD22 , ANI-1x, QM24 [58]
Table B.6: Protein Secondary Structure Codes. Summary of the standard single-letter codes used to represent protein secondary structure elements, as defined by the DSSP classification [20].

### v3

@@ 7 Interpreting MLFF representations
While MLFF atomistic features excel at representing local protein structure and chemistry, it remains unclear how they respond to structural perturbations or encode meaningful physical properties. To address this, the behavior of MLFF representations under specific conformational changes is examined. Such changes induce smooth, interpretable trajectories in the embedding space, suggesting that the features capture physical motions and respect locality, continuity, and symmetry.
Figure 6: Structure of the MACE embedding space for the rotating phenylalanine aromatic ring from Fig. 4. Left: a fragment of a protein with simulated rotation of the phenylalanine sidechain ring. Right: Mace embeddings of backbone CA atoms projected onto the first two principal components. The vertical axis separates the different rotation angles for visual clarity. Note that residues close to the ring ($50$ and $51$) exhibit a one-dimensional structure with nearly perfect $180$-degree periodicity. The farther residue $49$ shows some breakdown of this structure, while residue $40$, uninfluenced by the ring current, manifests a lack of relation between ring rotation and the embedding space structure.
@@
Rotated phenylalanine sidechain aromatic ring. Using the synthetically modified phenylalanine, we perform PCA on MACE embeddings obtained at every $\chi_{2}$ angle for the nearby CA atoms (residues $50$ and $51$ of PDB ID: 1ZV6). We observe that the first two principal components trace a smooth, one-dimensional curve, capturing the expected $180^{\circ}$ periodicity of the rotation. However, the effect diminishes with distance: residue $49$ begins to deviate from the periodic pattern, while distant residue $40$ exhibits no discernible structure, illustrating the spatial locality of the descriptor’s sensitivity.
Additional case studies further analyzing the embedding space for helix-to-strand unfolding MD trajectory and experiments related to inverting MLFF embeddings, are presented in Section  I and Section J.2, respectively, in the Appendix.
@@ 8 Discussion
In this work, we introduced a novel way to represent local protein environments by repurposing embeddings from MLFFs. To our knowledge, this is the first demonstration that MLFF latent spaces – trained exclusively on small-molecule quantum data – organize according to meaningful biochemical factors such as secondary structure, residue identity, protonation state, and chemical shift. We further show that these embeddings can be directly reused for diverse protein modeling tasks without retraining. This establishes MLFFs as general-purpose, reusable representation learners for structural biology and opens a new paradigm for leveraging pretrained molecular models in this domain.
@@
While our downstream models used frozen MLFF embeddings, task-specific fine-tuning of MLFFs is a promising direction that may yield further gains. Moreover, our chemical shift prediction for HA, while accurate, still slightly lags behind UCBShift, pointing to opportunities for future improvement. Notwithstanding these limitations, we believe that the proposed fully differentiable chemical shift predictor can be used to guide AlphaFold and similar generative models (Maddipatla et al., 2025) for determining structure from experimentally measured chemical shifts – an important task in protein NMR that has so far been hindered by the complexity of structure-to-chemical shift relation.
9 Ethics Statement
All authors have read and complied with the ICLR Code of Ethics. This work does not involve human subjects, personally identifiable information, or sensitive data. All datasets mentioned in Appendix E are publicly available, and we adhered to best practices to ensure privacy, fairness, and transparency. We do not foresee any direct ethical risks associated with the methods or results presented.
10 Reproducibility Statement
The code, trained models, and data-processing scripts are publicly available at https://github.com/mb012/MLFF_representation. In addition, complete details of the models and optimization parameters are provided in Appendix G.3. The hardware resources used to produce the results are specified in Appendix I.4. The loss functions, evaluation metrics, and details regarding ablation studies are specified in Appendix G. These details ensure that all results reported in the paper can be independently verified.
Acknowledgments
This work was supported by the Institute of Science and Technology Austria (ISTA) through the IPC grant “Generative Protein NMR” and by the Israeli Science Foundation (ISF) under grant number 1834/24. This research used resources of the Institute of Science and Technology Austria’s scientific computing cluster. S.V. was supported in part by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard.
Appendix A Summary of experiments
Table B.2: Experimental overview. Summary of key research questions, experimental design, baselines, main takeaways, and supporting evidence. “ESM family” includes embeddings from ESM2/ESMFold (Lin et al., 2023), ISM (Ouyang-Zhang et al., 2025), and ESM3 (Hayes et al., 2025). On the other hand, “learned” refers to the learned-embedding baseline described in Section 3.
Question Experiment Design Baselines Takeaway Evidence
1. Do MLFF representations capture structural and chemical properties? Train classifiers to predict amino acid identity (20 types) and secondary structure ($\alpha$-helix, $\beta$-strand, other) from MLFF embeddings. Visualize embedding space using UMAP. MACE, Egret, OrbNet, AIMNet, ESM family Egret best for amino acid classification; MACE/Egret best for secondary structure. UMAP shows clear clustering by chemical identity and structural motifs. Fig. 2
2. Can MLFF representations predict pKa values? Train regression models to predict pKa for ionizable residues (GLU, ASP, LYS, HIS) using MLFF embeddings. Target pKa values are approximated by PypKa’s Poisson-Boltzmann solver. PropKa, pKa-ANI, ESM family, learned, and MLFF variants AIMNet features achieve best performance. Table 1
3. Can MLFF embeddings detect distribution shifts and separate structural classes? (a) Measure likelihood of environments before/after Amber99 relaxation using kernel density estimation. (b) Construct conditional likelihoods for different secondary structures. Compare distributions before/after relaxation (a) Relaxed structures receive higher likelihoods. (b) Conditional likelihoods clearly separate secondary structure classes. Figs. A.7, A.11
4. Can MLFF-based predictors outperform state-of-the-art chemical shift prediction? Train GNNs on MLFF embeddings to predict NMR chemical shifts for backbone (N, CA, C, H, HA) and side-chain atoms. UCBShift2-X (SOTA), ESM family, learned MLFF-based predictors outperform UCBShift2-X for backbone and side-chain heavy atoms (except HA). MACE-based models perform best overall. Fig. 3, Table B.5, Table B.6
[…]
6. Can likelihood serve as an uncertainty estimate? Build unconditional likelihood model on training embeddings; correlate environment likelihood with chemical shift prediction error on test set. No direct baseline (novel metric) Low-likelihood environments consistently exhibit higher prediction errors, supporting use for uncertainty estimation. Fig. 5
7. How do MLFF embeddings respond to structural perturbations? Perform PCA on MACE embeddings during: (a) systematic phenylalanine ring rotation, (b) helix-to-strand MD simulation. Analyze embedding trajectories and spatial locality. Visual analysis of embedding trajectories Embeddings trace smooth trajectories reflecting structural changes, show expected periodicity, and reveal interpretable latent directions. Figs. 6, A.9, A.10
8. Can MLFF embeddings be inverted to recover structure? Optimize AlphaFold3 structures using only MACE descriptor matching as loss function. Test on alternate conformations in: 4OLE,3V3S,4NPU,5G51, 7EC8. Direct comparison to target conformations Backbone geometry recovered with high accuracy; side-chain orientations less precise. MLFF embeddings encode most but not all information needed for reconstruction. Fig. A.14,  A.15,  A.16
9. Comparative benchmarking and ablation studies (1) Benchmark all MLFF families on all tasks. (2) Ablate atom set selection. (3) Compare MLFF similarity metric to LoCoHD. MACE, Egret, OrbNet, AIMNet variants; LoCoHD metric MACE performs best on most tasks (AIMNet best for pKa). Full atom set improves performance; MLFF similarity outperforms LoCoHD. Tables B.3-B.11, Fig. A.13
Appendix B Additional Figures
Figure A.7: MACE-based likelihood captures structurally subtle distribution shifts in protein environments. The histogram of likelihoods of CA MACE representations of $26000$ protein environments curated from $225$ protein structures in the test set, before (red) and after (blue) relaxing the protein structures with an Amber99 force-field (Hornak et al., 2006).
@@ Appendix B Additional Figures
Figure A.9: Structure of the MACE embedding space for the unfolding helix from Fig. A.8. Depicted are Mace embeddings of the backbone CA atoms of residues $6$ and 77 projected onto the first two principal components. Observe an essentially one-dimensional structure of the trajectory in the PCA space.
[…]
@@
Figure A.12: Distribution of the secondary chemical shifts of the two stable conformers of 4OLE. Left: $100$ ns MD simulations (Rosenberg et al., 2024) of the two conformers (marked as A (containing a helix) and B (containing a linearly structured loop) according to the original PDB annotation). Lysine 63 is highlighted. Right: secondary shift distributions of lysine 63 CA (top) and CB (bottom) atoms over the MD trajectory. Note that the helical conformer exhibits an expected excess in the CA secondary shift and a defect in the CB secondary shift.
Figure A.13: Similarity of local protein environments using MACE and LoCoHD. In each plot, environments belonging to different class labels are represented as points while pairwise Euclidean distances approximate the dissimilarities as measured using the LoCoHD metric (Fazekas et al., 2024) and a metric between MACE likelihoods. Circles indicate each class variability as seen by the metric. Shown is conditioning by secondary structure (top row) and amino acid identity (bottom row). We conclude that MACE captures much better the similarities between related chemical and structural classes, even if LoCoHD has been explicitly designed for these tasks.
Figure A.14: AlphaFold3 guided structures recovered by inverting MACE descriptors. (A) The unguided AlphaFold3 structures (orange) of four proteins (3V3S, 4NPU, 5G51, and 7EC8) with experimentally determined alternate backbone conformations are shown alongside guided predictions (green), obtained by inverting the corresponding conformation’s MACE descriptor. All predictions are overlaid on the experimentally determined conformer (white) from the PDB (Burley et al., 2017). (B) The violin plots of the signed normalized distance distributions relative to conformations A (1-1) and B (11) in generated structures for the same cases, comparing unguided (orange) and descriptor-guided predictions (green / blue).
[…]
Figure A.16: Inverting MACE descriptors by optimizing atomic coordinates Shown in green is the structure of the protein 4OLE:B as predicted by AlphaFold3. The structure has an alternative conformation in the region 60 - 68, (shown in red is the crystallographic structure from the PDB), whose MACE environment descriptors were used to optimize the AlphaFold prediction. The solution of the optimization problem is depicted in pink.
Appendix C Quantitative Results & Tables
@@
ESM-Features + GCN MLFF-Features + GCN (Ours)
@@
Metric Class LocoHD ESM3 (Seq) ISM ESM3 (Struc) ESMFold Egret MACE OrbNet AIMNet
Precision [%] Weighted Mean $76.947$ $88.175$ $92.467$ $94.881$ $93.333$ $95.685$ $\mathbf{95.694}$ $92.500$ $94.716$
Mean $76.424$ $87.740$ $92.279$ $94.981$ $93.304$ $\mathbf{95.540}$ $95.528$ $92.233$ $94.522$
$\alpha$-helix $80.493$ $90.303$ $93.065$ $96.103$ $94.431$ $96.466$ $\mathbf{96.657}$ $93.900$ $95.762$
$\beta$-strand $75.623$ $91.124$ $95.555$ $95.252$ $95.757$ $\mathbf{96.248}$ $95.969$ $93.749$ $95.494$
Other $73.158$ $81.793$ $88.216$ $92.419$ $88.941$ $93.905$ $\mathbf{93.958}$ $89.049$ $92.309$
Recall [%] Weighted Mean $77.334$ $88.268$ $92.505$ $95.991$ $93.327$ $\mathbf{98.636}$ $95.717$ $92.574$ $94.752$
Mean $75.532$ $87.637$ $92.071$ $95.799$ $92.961$ $\mathbf{98.508}$ $95.355$ $91.977$ $94.335$
$\alpha$-helix $87.227$ $91.299$ $94.562$ $96.845$ $95.023$ $\mathbf{97.546}$ $97.522$ $95.548$ $96.805$
$\beta$-strand $89.035$ $92.859$ $95.818$ $96.792$ $96.338$ $\mathbf{97.967}$ $97.947$ $96.276$ $97.481$
Other $50.333$ $78.753$ $85.833$ $89.759$ $87.521$ $90.512$ $\mathbf{90.597}$ $84.106$ $88.720$
F1 [%] Weighted Mean $76.323$ $88.210$ $92.479$ $95.433$ $93.311$ $\mathbf{98.633}$ $95.693$ $92.509$ $94.720$
Mean $75.048$ $87.675$ $92.168$ $95.388$ $92.999$ $\mathbf{98.552}$ $95.427$ $92.073$ $94.412$
$\alpha$-helix $83.725$ $90.798$ $93.807$ $96.473$ $94.726$ $97.003$ $\mathbf{97.087}$ $94.717$ $96.281$
$\beta$-strand $81.783$ $91.983$ $95.687$ $96.016$ $96.047$ $\mathbf{97.100}$ $96.948$ $94.996$ $96.477$
Other $59.636$ $80.244$ $87.008$ $91.069$ $88.225$ $92.177$ $\mathbf{92.247}$ $86.507$ $90.479$
@@
Table B.4: Amino acid prediction accuracy. Reported are precision, recall, and F1 scores, each calculated with respect to the ground truth amino acid class. The best-performing model for each metric is indicated in bold. Unlike secondary structure prediction, per-class performance is not reported here due to the large number (2020) of amino acid types.
ESM-Features + GCN MLFF-Features + GCN (Ours)
@@
Metric Class LocoHD ESM3 (Seq) ISM ESM3 (Struc) ESMFold Egret MACE OrbNet AIMNet
Precision [%] Weighted Mean $97.794$ $91.081$ $97.919$ $96.864$ $98.884$ $\mathbf{99.129}$ $98.448$ $95.967$ $98.665$
Mean $98.012$ $89.923$ $96.891$ $96.828$ $97.821$ $\mathbf{98.758}$ $98.031$ $95.090$ $98.407$
Recall [%] Weighted Mean $97.805$ $91.089$ $96.929$ $96.964$ $98.642$ $\mathbf{99.133}$ $98.431$ $95.955$ $98.664$
Mean $97.699$ $89.484$ $95.922$ $96.803$ $97.823$ $\mathbf{98.932}$ $98.262$ $95.072$ $98.457$
F1 [%] Weighted Mean $97.785$ $91.079$ $97.421$ $96.914$ $98.763$ $\mathbf{99.130}$ $98.435$ $95.959$ $98.663$
Mean $97.839$ $89.686$ $96.404$ $96.815$ $97.822$ $\mathbf{98.842}$ $98.137$ $95.078$ $98.431$
@@
Table B.5: Backbone chemical shift prediction accuracy. Reported is the mean absolute error (MAE), in ppm, relative to the experimentally measured chemical shift values from RefDB (Zhang et al., 2003) for different atoms. The best-performing model for each metric is indicated in bold.
ESM features + GCN MLFF-features + GCN (Ours)
@@
Atom UCB Shift LocoHD Learned ESM2 ESM3 (Seq) ISM ESM3 (Struc) ESMFold MACE Egret Orb AIMNet
HA $\mathbf{0.165}$ N/A $0.405$ $0.247$ $0.235$ $0.220$ $0.193$ $0.203$ $0.180$ $0.176$ $0.200$ $0.201$
H $0.324$ N/A $0.528$ $0.402$ $0.393$ $0.377$ $0.600$ $0.365$ $0.300$ $\mathbf{0.295}$ $0.333$ $0.325$
CA $0.653$ $1.333$ $1.753$ $0.872$ $1.049$ $0.814$ $0.667$ $0.667$ $\mathbf{0.584}$ $0.599$ $0.683$ $0.660$
N $1.891$ $2.877$ $3.174$ $2.455$ $2.349$ $2.275$ $1.845$ $1.986$ $\mathbf{1.642}$ $1.667$ $2.015$ $1.875$
CB $0.758$ $1.140$ $1.326$ $0.923$ $1.092$ $0.884$ $0.758$ $0.775$ $0.716$ $\mathbf{0.705}$ $0.792$ $0.780$
C $0.827$ $1.297$ $1.517$ $1.011$ $1.038$ $0.923$ $0.842$ $0.876$ $0.744$ $\mathbf{0.743}$ $0.847$ $0.824$
Table B.6: Side chain chemical shift prediction accuracy. Reported is the mean absolute error (MAE), in ppm, relative to the experimentally measured chemical shift values from RefDB (Zhang et al., 2003) for different atoms. The best-performing model for each metric is indicated in bold.
MLFF-features + GCN (Ours)
Atom UCB Shift LocoHD MACE Egret Orb AIMNet
CG $0.595$ $0.647$ $0.567$ $\mathbf{0.563}$ $0.638$ $0.611$
CD $0.615$ $0.512$ $0.456$ $\mathbf{0.448}$ $0.521$ $0.482$
CD2 $1.116$ $1.195$ $1.035$ $1.085$ $1.201$ $1.749$
CG2 $0.773$ $0.833$ $\mathbf{0.713}$ $0.720$ $0.760$ $0.742$
CE $0.519$ $2.295$ $0.447$ $\mathbf{0.436}$ $0.451$ $0.446$
Table B.7: PropKa vs. MACE pKa comparison. Per-residue pKa prediction errors for Glutamic acid and Aspartic acid from PropKa’s training set. PropKa is evaluated in its native setting, using experimental pKa values derived from crystallographic PDB structures, whereas our MACE-based model was trained using simulated pKa values from AlphaFold2 structures. This comparison inherently favors PropKa, yet our model achieves lower error on more than half of the structures in PropKa’s training set.
Glutamic acid Aspartic acid
PDB ID PropKa Ours (MACE) PropKa Ours (MACE)
1PGA:A $0.413$ $\mathbf{0.294}$ $0.362$ $\mathbf{0.292}$
1IGD:A $0.489$ $\mathbf{0.182}$ $0.352$ $\mathbf{0.274}$
1A2P:A $\mathbf{0.525}$ $1.046$ $0.426$ $\mathbf{0.403}$
4ICB:A $0.579$ $\mathbf{0.546}$ $\mathbf{0.299}$ $1.018$
1BEO:A - - $0.730$ $\mathbf{0.511}$
4LZT:A $\mathbf{0.629}$ $0.665$ $0.488$ $\mathbf{0.404}$
3RN3:A $0.548$ $\mathbf{0.462}$ $0.423$ $\mathbf{0.412}$
2RN2:A $0.274$ $\mathbf{0.185}$ $0.683$ $\mathbf{0.585}$
2OVO:A $0.377$ $\mathbf{0.376}$ $\mathbf{0.180}$ $1.571$
1XNB:A $\mathbf{1.001}$ $1.270$ $\mathbf{0.436}$ $1.020$
135L:A $0.470$ $1.700$ $\mathbf{0.480}$ $1.043$
Table B.8: pKa-ANI vs. MACE pKa performance. Mean absolute pKa prediction error (MAE) for four amino acids using pKa-ANI and our MACE-based model, evaluated on the PKAD-R dataset (Ancona et al., 2023), a curated version of the dataset on which pKa-ANI was originally trained and tested. While pKa-ANI is evaluated in its native setting, our model was trained on simulated pKa values derived from AlphaFold2 structures, making this comparison inherently favorable to pKa-ANI. Nevertheless, our model outperforms pKa-ANI on three of the four amino acids; for Histidine, we hypothesize that the poor performance is due to an insufficient number of training examples for that residue type.
Amino acid pKa-ANI Ours (MACE)
Glutamic Acid $0.696$ $\mathbf{0.695}$
Aspartic Acid $0.978$ $\mathbf{0.831}$
Lysine $0.648$ $\mathbf{0.360}$
Histidine $\mathbf{0.899}$ $0.925$
Table B.9: Environment radius ablation. Prediction error (ppm; lower is better) for backbone chemical shift prediction as a function of the local environment radius $r$. We compare $r=4,5,6$ Å using MACE embeddings as input to the GCN. The best-performing radius (55 Å) is highlighted in bold.
Atom 44Å 55Å 66Å
HA $0.194$ $\mathbf{0.180}$ $0.191$
H $0.320$ $\mathbf{0.300}$ $0.318$
CA $0.620$ $\mathbf{0.584}$ $0.627$
N $1.779$ $\mathbf{1.642}$ $1.768$
CB $0.758$ $\mathbf{0.716}$ $0.754$
C $0.798$ $\mathbf{0.744}$ $0.804$
Table B.10: MLFF layer ablation. Per-atom mean absolute error (MAE) in backbone chemical-shift prediction for different choices of descriptor layer. For MACE, L1 and L2 denote descriptors taken from the first and second message-passing layers, and L1+2 denotes their concatenation. For Orb, L5, L10, and L15 denote descriptors taken from the 5th, 10th, and 15th message-passing layers, respectively. For each atom type and model, the best-performing layer configuration is highlighted in bold.
Atom MACE L1 MACE L2 MACE L1+2 Orb L5 Orb L10 Orb L15
C 0.797 0.755 0.744 0.836 0.836 0.847
CA 0.619 0.593 0.584 0.678 0.660 0.683
CB 0.756 0.705 0.716 0.778 0.777 0.792
H 0.318 0.300 0.300 0.330 0.331 0.333
HA 0.191 0.178 0.180 0.198 0.196 0.200
N 1.771 1.666 1.642 1.929 1.889 2.015
@@
Table B.11: Machine Learning Force Field (MLFF) metadata. Listed are the model type, dimensionality, geometric properties of the representations, and training dataset for MLFFs in Section 2.
@@
MACE large $448$ $S_{n}$ & $O(3)$ invariant QM9 (Ramakrishnan et al., 2014), MD22, ANI-1x
OrbNet-v2 $256$ $S_{n}$ & $E(3)$ invariant QM9, GDB13 (Cheng et al., 2019), DrugBank (Law et al., 2014)
AIMNet2 $256$ $S_{n}$ & $E(3)$ invariant ANI-1x (Smith et al., 2018)
Egret1 $384$ $S_{n}$ & $O(3)$ invariant QM9, MD22 (Chmiela et al., 2023) , ANI-1x, QM24 (Ruddigkeit et al., 2012)
Table B.12: Protein Secondary Structure Codes. Summary of the standard single-letter codes used to represent protein secondary structure elements, as defined by the DSSP classification (Frishman & Argos, 1995).